If 3 times a certain number is increased by 4 the result is 28 If x is "the number," which of the following equations could be used to solve for the number?
a.3x + 4 = 28
b.3(x + 4) = 28
c.3x = 4 + 28


Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(10682)

 
Hi,
If 3 times a certain number (x) is increased by 4 the result is 28
Write as we Read
3x + 4 = 28

Answer by stanbon(57384)

If 3 times a certain number is increased by 4 the result is 28 If x is "the number," which of the following equations could be used to solve for the number?
a.3x + 4 = 28
b.3(x + 4) = 28
c.3x = 4 + 28
---
Ans: a
------------
Cheers,
Stan H.